OpenStudy (anonymous):

Simplify 7(square root of 5)-1(square root of 5)

OpenStudy (anonymous):

These are fun. Think of the square roots as variables. For example: 7x -1x = ? Same thing here \[7\sqrt{5} - 1\sqrt{5} =?\] give it a shot :D

OpenStudy (anonymous):

6root5